OpenStudy (anonymous):

solve the exponential equation using the method of relating the bases by first rewriting the equation in the form of e^u=e^v 1/e^6x=sqrootof e/e^3-x

OpenStudy (anonymous):

I cannot make my calculator do this for some reason

OpenStudy (vijay):

u mean 1/e^6x=sqrootof (e/e^3-x)

OpenStudy (anonymous):

\[1/e ^{6x}=\sqrt{e}/e ^{x-3}\]

OpenStudy (anonymous):

power of 3-x - sorry

OpenStudy (anonymous):

-6x=1/2-x+3 x=-.7

OpenStudy (anonymous):

it says -.7 is not right

OpenStudy (anonymous):

and gave me a new problem, grrrrr